# Hardware Acceleration for Neural Networks: A Comprehensive Survey

### v1

@@ 2.5.4. FPGA-specific utilization challenges
FPGA utilization depends on balancing pipelines, memory ports, and limited DSP/BRAM resources. Designs can become bandwidth-bound or routing-limited, and achieving high utilization across diverse models is difficult without reconfiguration or overprovisioning (Nurvitadhi et al., 2017; Venieris and Bouganis, 2017). For attention/Transformer workloads, limited on-chip buffering can reduce utilization by forcing frequent external memory access (Zhang et al., 2024a; Vaswani et al., 2017). In cloud inference services, utilization is also shaped by batching and multi-tenant scheduling, which may be dominated by service-level constraints rather than by the FPGA datapath (Fowers et al., 2018). Toolchain maturity affects achievable utilization because memory partitioning and pipelining decisions are often compiler/HLS-limited (Venieris and Bouganis, 2017).
@@ 2.5.7. Neuromorphic utilization challenges
Neuromorphic utilization depends on spike sparsity and event locality. When activity is dense, communication contention increases and the advantage of event-driven computation diminishes (Davies et al., 2018; Merolla et al., 2014). Utilization can also be limited by the mapping of the network to the hardware fabric; poor mapping can concentrate spikes and create hotspots even when global activity is low (Davies et al., 2018). Converting conventional models to spiking forms can reduce effective sparsity and therefore reduce utilization benefits (Gale et al., 2019). Consequently, utilization must be evaluated jointly with the encoding, network design, and application workload (Coleman et al., 2019).
@@ 2.6. Benchmarking and reproducibility
Figure 8. Benchmarking and reproducibility challenges
Benchmark suites such as MLPerf aim to standardize tasks, accuracy targets, and reporting rules for both training and inference, improving comparability across vendors and platforms (Mattson et al., 2020). However, representative evaluation still requires careful choice of workload mix: production systems often run diverse models, experience non-stationary request distributions, and care about tail latency and cost, not only peak throughput. End-to-end benchmarks and competitions (e.g., time-to-train style evaluations) complement kernel microbenchmarks by capturing the interaction between the model, the input pipeline, the optimizer, and the distributed system (Coleman et al., 2019).
@@ 2.6.4. FPGA benchmarking challenges
FPGA results are sensitive to toolchain choices (HLS vs. RTL), frequency targets, and memory configuration, and they often depend on end-to-end pipeline integration (host I/O, streaming, reconfiguration) (Nurvitadhi et al., 2017; Fowers et al., 2018). Reporting should include resource utilization (DSP/BRAM/LUT), achieved clock, and I/O constraints, because these determine whether the design is compute- or bandwidth-limited (Venieris and Bouganis, 2017; Williams et al., 2009). Transformer benchmarking should include sequence lengths and attention implementation details, since buffering and external memory access dominate performance (Zhang et al., 2024a; Vaswani et al., 2017). In cloud settings, multi-tenant effects and batching policies can change both throughput and tail latency and should be reported explicitly (Fowers et al., 2018).
@@ 2.6.7. Neuromorphic benchmarking challenges
Neuromorphic benchmarks must report spike encoding, activity sparsity, and application-level accuracy, since performance and energy benefits depend strongly on event sparsity and representation (Davies et al., 2018; Merolla et al., 2014). Benchmarks should also report event rates and communication patterns, because routing congestion can dominate latency and energy when activity increases (Davies et al., 2018). Comparing against dense accelerators requires careful definition of equivalent workloads and latency/accuracy targets, and should account for preprocessing and encoding overheads (Coleman et al., 2019; Gale et al., 2019). As with analog accelerators, end-to-end evaluation is essential: reporting only chip-level energy without the full pipeline can be misleading (Mattson et al., 2020).
@@ 3. Hardware Accelerator Architectures for Neural Networks
Neural-network acceleration has evolved into a heterogeneous ecosystem where different platforms optimize different points in the design space. At a high level, accelerators implement dense linear algebra efficiently (GEMM/conv), reduce data movement through carefully designed memory hierarchies and dataflows, and increasingly incorporate support for sparsity, mixed precision, and attention-centric kernels. This section surveys major architectural families and discusses how they are used for inference and training across common model classes (CNNs, RNNs, GNNs, and Transformers/LLMs) (Williams et al., 2009; Vaswani et al., 2017; Jouppi et al., 2017).
Figure 9. Neural-network acceleration
Figure 9 provides an overview of the neural-network acceleration pipeline and where architectural choices impact performance.
@@ 3.1. GPU architectures
GPUs remain the dominant platform in many production and research settings due to their programmability, mature software ecosystem, and strong support for both training and inference. Architecturally, modern GPUs combine wide SIMT execution, deep multithreading to hide latency, and high-throughput memory systems (large on-chip caches and register files, plus high-bandwidth off-chip memory), together with specialized matrix-multiply units (tensor cores) that accelerate dense linear algebra at reduced precision (Williams et al., 2009; Micikevicius et al., 2018). In practice, GPU performance is inseparable from the software stack: operator libraries (e.g., highly tuned convolution and GEMM kernels), graph compilers, and kernel DSLs determine whether a model saturates compute or becomes bottlenecked by memory bandwidth, kernel launch overheads, or synchronization (Chetlur et al., 2014; Chen et al., 2018; Tillet et al., 2019).
@@ 3.1.2. Training on GPUs
Training emphasizes throughput, numerical stability, and scalability. Mixed-precision training has become a standard approach to improve performance while maintaining convergence, typically using FP16/BF16 compute with FP32 accumulation and loss-scaling or similar techniques (Micikevicius et al., 2018). Memory footprint is a dominant constraint: activations, gradients, and optimizer states can exceed device memory for large models, motivating activation checkpointing, optimizer-state partitioning, and distributed parallelism strategies (data, tensor, pipeline, and expert parallelism) (Rajbhandari et al., 2020; Shoeybi et al., 2019; Shazeer et al., 2017). These strategies trade off communication volume, synchronization, and pipeline bubbles against memory savings and compute efficiency.
@@ 3.2. Domain-specific tensor processors (TPUs and NPUs)
Figure 10. TPU architecture
Figure 10 illustrates a TPU-style tensor-processor organization used to motivate compilation, tiling, and on-chip reuse. The “tensor processor” category includes both data center-class training systems (e.g., TPU pods with specialized interconnect) and embedded inference engines. Across both, the core architectural challenge is balancing compute with on-chip reuse: increasing MAC throughput is only beneficial if activations and weights can be delivered at sufficient bandwidth without thrashing the memory hierarchy. This is especially visible for attention-heavy models, where KV-cache and intermediate activations introduce large, streaming memory demands that may not match the accelerator’s preferred dense reuse patterns (Vaswani et al., 2017; Kwon et al., 2023).
@@ 3.2.1. Inference on TPUs/NPUs
Inference on tensor processors benefits from predictable dense kernels and static shapes, enabling aggressive compilation, operator fusion, and on-chip reuse. Many deployments target integer arithmetic (INT8 and below) to reduce energy and improve throughput, and hardware often includes dedicated datapaths for quantized GEMM/conv and activation functions to avoid expensive conversions (Jacob et al., 2018). For CNNs, static-shape inference allows compilers to pre-plan tiling and buffering to maximize weight/activation reuse; for RNNs, throughput is more sensitive to sequence length and batch size due to limited parallelism and recurrent dependencies.
@@ 3.3.1. Inference on ASICs
Figure 11. Challenge for ASIC training
@@ 3.4. FPGA-based accelerators
FPGAs offer reconfigurability and the ability to tailor datapaths, memory access, and I/O interfaces to a workload, making them attractive for edge inference, low-latency pipelines, and rapid architectural experimentation. FPGA accelerators commonly use deeply pipelined streaming architectures, custom precision, and specialized memory layouts, and can integrate tightly with sensors and networking to reduce end-to-end latency (Nurvitadhi et al., 2017; Umuroglu et al., 2017; Venieris and Bouganis, 2017; Xu et al., 2025d, b, a). Unlike GPUs, FPGAs allow designers to specialize the datapath for a fixed model (or model family) and to exploit determinism in the workload, which can yield strong energy efficiency when memory bandwidth is sufficient and the operator set is well covered.
The FPGA design space also includes cloud-scale deployments where reconfigurable logic is used as a shared inference service. In these settings, throughput and tail latency depend not only on compute pipelines but also on host-device interfaces, batching policies, and the ability to switch between models with minimal reconfiguration overhead (Fowers et al., 2018; Nurvitadhi et al., 2017). Toolchains and abstraction layers (HLS, OpenCL, compiler-based mapping) play an outsized role in productivity and performance portability, influencing how quickly designs can adapt to new model operators and sparsity patterns (Venieris and Bouganis, 2017).
Figure 12. FPGA Architecture
Figure 12 sketches a representative FPGA-based acceleration setup used to discuss streaming datapaths, buffering, and I/O constraints.

### v3

@@ 2.5.2. TPU/NPU-specific utilization challenges
The rigidity of systolic arrays poses a utilization risk for non-GEMM operations. While matmuls run at near-peak efficiency, element-wise ops (activations, normalization) or reductions often run on separate vector units with much lower throughput. If a model is heavy on LayerNorm or Softmax (relative to matmul), these vector units become the bottleneck, leaving the massive systolic arrays idle. This “Amdahl’s Law” effect means that as matrix engines get faster, the utilization bottleneck shifts aggressively to the vector/transcendental units and data-reshaping logic (Jouppi et al., 2017; Williams et al., 2009).
@@ 2.5.3. ASIC-specific utilization challenges
A subtle utilization killer for ASICs is the “batching mismatch.” If an ASIC is designed with a very wide vector width (e.g., 512-byte SIMD) to maximize peak TOPS, it requires a minimum batch size to fill those vectors. In real-time serving where batch size might be 1, the hardware effectively runs at a fraction of its capacity (e.g., 1/32th utilization). Techniques like “batch-1 optimization” or systolic arrays that stream weights (weight-stationary) can help, but they often require different on-chip buffering strategies than training-optimized designs. Thus, an ASIC built for training may show abysmal utilization in inference, and vice-versa (Hennessy and Patterson, 2019).
@@ 2.5.4. FPGA-specific utilization challenges
FPGA utilization depends on balancing pipelines, memory ports, and limited DSP/BRAM resources. Designs can become bandwidth-bound or routing-limited, and achieving high utilization across diverse models is difficult without reconfiguration or overprovisioning (Nurvitadhi et al., 2017; Venieris and Bouganis, 2016). For attention/Transformer workloads, limited on-chip buffering can reduce utilization by forcing frequent external memory access (Zhang et al., 2024a; Vaswani et al., 2017). In cloud inference services, utilization is also shaped by batching and multi-tenant scheduling, which may be dominated by service-level constraints rather than by the FPGA datapath (Fowers et al., 2018). Toolchain maturity affects achievable utilization because memory partitioning and pipelining decisions are often compiler/HLS-limited (Venieris and Bouganis, 2016).
Furthermore, “logic utilization” on FPGAs has a practical ceiling (often around 70-80%) before routing congestion makes timing closure impossible. A design that theoretically uses 95% of the DSPs might fail to route or run at a very low clock frequency, effectively reducing throughput. This routing overhead means that the “usable peak” is significantly lower than the “datasheet peak.” Moreover, the time spent reconfiguring the FPGA (programming the bitstream) is dead time; in multi-tenant clouds, this reconfiguration latency discourages fine-grained time-sharing, leading to static partitioning that may leave regions underutilized (Fowers et al., 2018; Williams et al., 2009).
@@ 2.5.5. LPU/LLM-serving utilization challenges
Specifically, the “prefill-decode imbalance” creates utilization gaps. During prefill, compute units are saturated; during decode, they are starved by memory bandwidth. In a naive system, the compute logic sits 90% idle during the long decode phase. Mixed-phase scheduling (running prefill for request A while request B is in decode) can recover some utilization, but it requires sophisticated resource isolation to prevent the prefill burst from destroying the decode latency Service Level Objective (SLO). LPU architectures often include specialized scheduling hardware to manage this mixing without Operating System (OS) overhead, aiming to keep the matrix units fed even when request phases diverge (Groq, 2024; Kwon et al., 2023).
@@ 2.5.6. In-/near-memory and analog utilization challenges
The “mapping fragmentation” problem is severe for crossbars. If a layer’s weight matrix doesn’t perfectly align with the physical crossbar size (e.g., a 100x100 matrix on a 128x128 array), the unused cells (rows/columns) are wasted area and power. Unlike digital memories that can be packed, analog weights are spatially fixed. This leads to low effective utilization for layers with odd shapes or for the tail ends of large matrices. Solutions involve virtualizing the arrays or using complex interconnects to route inputs to sub-tiles, but these add overhead. Consequently, utilization is highest for uniform, large-model layers and drops for irregular, optimized architectures (like MobileNets) (Shafiee et al., 2016; Chi et al., 2016).
@@ 2.5.7. Neuromorphic utilization challenges
Neuromorphic utilization depends on spike sparsity and event locality. When activity is dense, communication contention increases and the advantage of event-driven computation diminishes (Davies et al., 2018; Merolla et al., 2014). Utilization can also be limited by the mapping of the network to the hardware fabric; poor mapping can concentrate spikes and create hotspots even when global activity is low (Davies et al., 2018). Converting conventional models to spiking forms can reduce effective sparsity and therefore reduce utilization benefits (Gale et al., 2019). Consequently, utilization must be evaluated jointly with the encoding, network design, and application workload (Coleman et al., 2017).
Moreover, “temporal utilization” in neuromorphic chips can be low if the input events are bursty. The hardware must be provisioned for the peak event rate to avoid dropping spikes, but during quiet intervals, the dedicated routing and update logic sits idle. Unlike clock-gated digital logic where idle means zero dynamic power, neuromorphic circuits often have leakage or bias currents that persist. Thus, the effective utilization (useful work per unit time / provisioned capacity) can be quite low for sporadic signals, challenging the TCO justification unless the standby power is exceptionally well managed (Davies et al., 2018).
@@ 2.6. Benchmarking and reproducibility
Figure 8. Benchmarking and reproducibility challenges in hardware acceleration, underscoring the sensitivity of results to software versions, compiler settings, model variants, and measurement methodology.
Benchmark suites such as MLPerf aim to standardize tasks, accuracy targets, and reporting rules for both training and inference, improving comparability across vendors and platforms (Mattson et al., 2020). However, representative evaluation still requires careful choice of workload mix: production systems often run diverse models, experience non-stationary request distributions, and care about tail latency and cost, not only peak throughput. End-to-end benchmarks and competitions (e.g., time-to-train style evaluations) complement kernel microbenchmarks by capturing the interaction between the model, the input pipeline, the optimizer, and the distributed system (Coleman et al., 2017).
@@ 2.6.1. GPU benchmarking challenges
Furthermore, benchmarking often ignores the “warm-up” phase, Just-In-Time (JIT) compilation time, and memory fragmentation effects, which are critical in production. A benchmark that measures steady-state throughput after 1000 iterations hides the initial 30-second compilation delay or the gradual performance degradation from heap fragmentation. For dynamic serving, measuring sustainable throughput (throughput at a fixed latency SLO) is more honest than peak throughput, but harder to standardize. Benchmarks should therefore report the “latency-throughput curve” rather than a single point, capturing the behavior under load (Mattson et al., 2020; Kwon et al., 2023).
@@ 2.6.2. TPU/NPU benchmarking challenges
The closed nature of some NPU software stacks adds a “black box” variability. An OS update or driver patch might change the NPU frequency governor or memory allocation policy, altering results significantly. Unlike open GPU kernels (e.g., in Triton or CUDA), NPU behavior is often hidden behind firmware blobs. This makes attribution difficult: did performance improve because of the hardware or because the compiler recognized a specific subgraph pattern? Fair benchmarking requires freezing not just the model but the entire firmware/driver image (Mazumder et al., 2021).
@@ 2.6.3. ASIC benchmarking challenges
Additionally, ASIC power measurements are often reported at the “chip” level, excluding DRAM or host power. This can be misleading for memory-bound workloads where DRAM power is 30-50% of the total. A “system-level” benchmark (wall power) is necessary to compare an ASIC against a GPU fairly. Also, many research ASICs report simulation results rather than silicon measurements; simulation assumptions (e.g., zero DRAM latency, perfect clock gating) should be scrutinized. Benchmarks should explicitly state whether numbers are measured on silicon, emulated on FPGA, or estimated via architectural simulators (Hennessy and Patterson, 2019; Williams et al., 2009).
@@ 2.6.4. FPGA benchmarking challenges
FPGA results are sensitive to toolchain choices (HLS vs. RTL), frequency targets, and memory configuration, and they often depend on end-to-end pipeline integration (host I/O, streaming, reconfiguration) (Nurvitadhi et al., 2017; Fowers et al., 2018). Reporting should include resource utilization (DSP/BRAM/LUT), achieved clock, and I/O constraints, because these determine whether the design is compute- or bandwidth-limited (Venieris and Bouganis, 2016; Williams et al., 2009). Transformer benchmarking should include sequence lengths and attention implementation details, since buffering and external memory access dominate performance (Zhang et al., 2024a; Vaswani et al., 2017). In cloud settings, multi-tenant effects and batching policies can change both throughput and tail latency and should be reported explicitly (Fowers et al., 2018).
A common pitfall is benchmarking only the “kernel execution time” on the FPGA, ignoring the PCIe transfer time to move data to/from the device. For small batches, this transfer dominates. Benchmarks must report end-to-end latency including all data movement. Also, precision matters: comparing an 4-bit Integer (INT4) FPGA implementation against an 16-bit Floating Point (FP16) GPU implementation is an apples-to-oranges comparison unless accuracy metrics are also provided. The “Pareto frontier” of Accuracy vs. Throughput/Watt is the only robust way to compare flexible precision FPGAs against fixed-precision logic (Nurvitadhi et al., 2017; Umuroglu et al., 2017).
@@ 2.6.5. LPU/LLM-serving benchmarking challenges
Furthermore, the choice of “random” vs. “real” text inputs affects benchmarking for LLMs due to tokenization and sparsity. If a benchmark uses random noise, the tokenizer might produce a different number of tokens than for English text, skewing throughput numbers. If the model uses content-dependent sparsity (MoE or early exit), random data might trigger worst-case paths. Therefore, standard datasets (like ShareGPT or Alpaca) are preferred over synthetic tensors to capture realistic execution divergence and cache hit rates (Mattson et al., 2020; Kwon et al., 2023).
@@ 2.6.6. In-/near-memory and analog benchmarking challenges
The definition of “operation” in analog is tricky. Is a noisy, low-precision MAC equivalent to a deterministic digital MAC? Benchmarks should use “effective” throughput at a target accuracy (e.g., ImageNet Top-1) rather than raw TOPS. If the analog chip requires a larger model or retraining to regain accuracy lost to noise, that overhead must be penalized. A metric like “ISO-accuracy energy efficiency” is more meaningful than “peak efficiency at unknown accuracy” (Chi et al., 2016; Jacob et al., 2018).
@@ 2.6.7. Neuromorphic benchmarking challenges
Neuromorphic benchmarks must report spike encoding, activity sparsity, and application-level accuracy, since performance and energy benefits depend strongly on event sparsity and representation (Davies et al., 2018; Merolla et al., 2014). Benchmarks should also report event rates and communication patterns, because routing congestion can dominate latency and energy when activity increases (Davies et al., 2018). Comparing against dense accelerators requires careful definition of equivalent workloads and latency/accuracy targets, and should account for preprocessing and encoding overheads (Coleman et al., 2017; Gale et al., 2019). As with analog accelerators, end-to-end evaluation is essential: reporting only chip-level energy without the full pipeline can be misleading (Mattson et al., 2020).
Finally, the “baseline” problem is pervasive: neuromorphic papers often compare against unoptimized CPU/GPU code. A fair comparison requires comparing against a highly optimized sparse-dense matrix engine or a dedicated low-power edge NPU running a quantized version of the same model. Benchmarking suites like “NeuroBench” attempt to standardize this by defining tasks where temporal dynamics are essential, ensuring that the event-driven nature is actually exercising a relevant capability rather than just running MNIST inefficiently (Davies et al., 2018; Coleman et al., 2017).
@@ 3. Hardware Accelerator Architectures for Neural Networks
Neural-network acceleration has evolved into a heterogeneous ecosystem where different platforms optimize different points in the design space. At a high level, accelerators implement dense linear algebra efficiently (GEMM/conv), reduce data movement through carefully designed memory hierarchies and dataflows, and increasingly incorporate support for sparsity, mixed precision, and attention-centric kernels. This section surveys major architectural families and discusses how they are used for inference and training across common model classes (CNNs, RNNs, GNNs, and Transformers/LLMs) (Williams et al., 2009; Vaswani et al., 2017; Jouppi et al., 2017). Figure 9 provides an overview of the neural-network acceleration pipeline and where architectural choices impact performance.
Figure 9. Overview of the neural-network acceleration pipeline, mapping the flow from high-level model definitions through graph compilation and optimization to execution on specialized hardware architectures.
@@ 3.1. GPU architectures
GPUs remain the dominant platform in many production and research settings due to their programmability, mature software ecosystem, and strong support for both training and inference. Architecturally, modern GPUs combine wide Single Instruction, Multiple Threads (SIMT) execution, deep multithreading to hide latency, and high-throughput memory systems (large on-chip caches and register files, plus high-bandwidth off-chip memory), together with specialized matrix-multiply units (tensor cores) that accelerate dense linear algebra at reduced precision (Williams et al., 2009; Micikevicius et al., 2018). In practice, GPU performance is inseparable from the software stack: operator libraries (e.g., highly tuned convolution and GEMM kernels), graph compilers, and kernel DSLs determine whether a model saturates compute or becomes bottlenecked by memory bandwidth, kernel launch overheads, or synchronization (Chetlur et al., 2014; Chen et al., 2018; Tillet et al., 2019).
@@ 3.1.2. Training on GPUs
Training emphasizes throughput, numerical stability, and scalability. Mixed-precision training has become a standard approach to improve performance while maintaining convergence, typically using 16-bit Floating Point (FP16)/Brain Floating Point (BF16) compute with 32-bit Floating Point (FP32) accumulation and loss-scaling or similar techniques (Micikevicius et al., 2018). Memory footprint is a dominant constraint: activations, gradients, and optimizer states can exceed device memory for large models, motivating activation checkpointing, optimizer-state partitioning, and distributed parallelism strategies (data, tensor, pipeline, and expert parallelism) (Rajbhandari et al., 2020; Shoeybi et al., 2019; Shazeer et al., 2017). These strategies trade off communication volume, synchronization, and pipeline bubbles against memory savings and compute efficiency.
@@ 3.2. Domain-specific tensor processors (TPUs and NPUs)
Figure 10. Representative TPU architecture featuring a large systolic array for dense matrix multiplication, coupled with unified on-chip buffers and a specialized control unit to maximize data reuse.
@@ 3.2.1. Inference on TPUs/NPUs
Figure 10 illustrates a TPU-style tensor-processor organization used to motivate compilation, tiling, and on-chip reuse. The “tensor processor” category includes both data center-class training systems (e.g., TPU pods with specialized interconnect) and embedded inference engines. Across both, the core architectural challenge is balancing compute with on-chip reuse: increasing MAC throughput is only beneficial if activations and weights can be delivered at sufficient bandwidth without thrashing the memory hierarchy. This is especially visible for attention-heavy models, where KV-cache and intermediate activations introduce large, streaming memory demands that may not match the accelerator’s preferred dense reuse patterns (Vaswani et al., 2017; Kwon et al., 2023). Inference on tensor processors benefits from predictable dense kernels and static shapes, enabling aggressive compilation, operator fusion, and on-chip reuse. Many deployments target integer arithmetic (8-bit Integer (INT8) and below) to reduce energy and improve throughput, and hardware often includes dedicated datapaths for quantized GEMM/conv and activation functions to avoid expensive conversions (Jacob et al., 2018). For CNNs, static-shape inference allows compilers to pre-plan tiling and buffering to maximize weight/activation reuse; for RNNs, throughput is more sensitive to sequence length and batch size due to limited parallelism and recurrent dependencies.
@@ 3.3.1. Inference on ASICs
Figure 11. Challenges in designing ASICs for training, highlighting the need for higher precision, backward-pass support, and flexibility to handle evolving optimization algorithms compared to inference-only designs.
@@ 3.4. FPGA-based accelerators
FPGAs offer reconfigurability and the ability to tailor datapaths, memory access, and I/O interfaces to a workload, making them attractive for edge inference, low-latency pipelines, and rapid architectural experimentation. FPGA accelerators commonly use deeply pipelined streaming architectures, custom precision, and specialized memory layouts, and can integrate tightly with sensors and networking to reduce end-to-end latency (Nurvitadhi et al., 2017; Umuroglu et al., 2017; Venieris and Bouganis, 2016). Unlike GPUs, FPGAs allow designers to specialize the datapath for a fixed model (or model family) and to exploit determinism in the workload, which can yield strong energy efficiency when memory bandwidth is sufficient and the operator set is well covered.
The FPGA design space also includes cloud-scale deployments where reconfigurable logic is used as a shared inference service. In these settings, throughput and tail latency depend not only on compute pipelines but also on host-device interfaces, batching policies, and the ability to switch between models with minimal reconfiguration overhead (Fowers et al., 2018; Nurvitadhi et al., 2017). Toolchains and abstraction layers (HLS, OpenCL, compiler-based mapping) play an outsized role in productivity and performance portability, influencing how quickly designs can adapt to new model operators and sparsity patterns (Venieris and Bouganis, 2016).
Figure 12. FPGA-based acceleration architecture illustrating a streaming datapath approach with custom buffering and I/O integration, designed to minimize off-chip memory access for deterministic inference.